OpenStudy (anonymous):

help please,how do you solve 2^x=10 ??? help please,how do you solve 2^x=10 ??? @Mathematics

OpenStudy (anonymous):

\[x=\frac{\log(10)}{\log(2)}\]

OpenStudy (ash2326):

take log both sides xlog2 =log 10 assuming base to be 10 x*0.414=1 now find x =1/0.414

OpenStudy (anonymous):

or if you prefer \[x=\frac{1}{\log(2)}\] or even \[\frac{\ln(10)}{\ln(2)}\]

OpenStudy (anonymous):

solve \[b^x=A\] for x via \[x=\frac{\log(A)}{\log(b)}\] this is called the "change of base" formula

OpenStudy (anonymous):

ok htank u all :)